OpenStudy (anonymous):

C = 3h - ax a. solve for "h" b. solve for "a" c. solve for "x" I think I got (a.)- C = 3h - ax C + ax = 3h C + ax ------ = h 3 -but for (b.) and (c.), all I've got is: -3h + C = -ax How do I get "a" away from "x", on either problem?

OpenStudy (precal):

part a is correct

OpenStudy (precal):

for part b divide both sides by -a recall -ax is -a times x the opposite of times is divide

OpenStudy (precal):

sorry my comment is for part c divide both sides by -a for part b divide both sides by x

OpenStudy (anonymous):

So, in the end: (a.) C + ax ------- = h 3 (b.) -3h + C -------- = a -x (c.) -3h + C -------- = x -a Correct?

OpenStudy (precal):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (precal):

Anytime :)